OpenStudy (anonymous):

Find the slope of the following line. The line 5 x - 3 y = 4 m = Find the slope of the following line. The line 5 x - 3 y = 4 m = @Mathematics

OpenStudy (anonymous):

5x-3y = 4 3y= 5x-4 y =5x-4/3 ( y=mx+b ) -- > m=5/3